Find a direct relationship between x and y. x= t-6 y= t^2-5t
Answer by jim_thompson5910(35256)

x= t-6

x+6 = t

t = x+6

----------------------------------------------------------------

y= t^2-5t

y = (x+6)^2-5(x+6)

y = x^2+12x+36-5x-30

y = x^2+7x+6

So the relationship is y = x^2+7x+6

note: you can factor to get y = (x+1)(x+6)